Let f(x) be the function 2 x - 2. Then the quotient
\frac{f(3+h)-f(3)}{h} can be simplified to

Answer by richard1234(7193)

Find f(3+h) and f(3) first.

f(3+h) = 2(3+h) - 2 = 2h + 4
f(3) = 2(3) - 2 = 4

Then,